# DiffMS: Diffusion Generation of Molecules Conditioned on Mass Spectra

### v1

@@ 1 Introduction
In this paper, we focus on the “inverse” MS problem and develop a novel machine learning framework for chemical structure generation given a mass spectrum, sometimes also described as de novo generation. A recent study of de novo generation from MS shows that all of the methods tested suffer from near-zero structural accuracy (Bushuiev et al., 2024). Among prior approaches to this task are language models that are trained to convert tokenized m/z values and intensities as inputs to SMILES strings as outputs (Butler et al., 2023; Litsa et al., 2023); however, these autoregressive language models do not capture the permutation-invariant nature of mass spectra and molecules, nor do they incorporate chemical formula constraints as helpful prior knowledge. Another family of approaches utilizes intermediate representations such as scaffolds (Wang et al., 2025) or fingerprints (Stravs et al., 2022) before generating chemical structures, which are arguably more chemically interpretable and leverage additional amounts of structure-only data available, but have not necessarily led to significant performance improvement on benchmarks. Compared to the complete structural elucidation challenge of the “inverse” MS problem, the chemical formula of the unknown molecule is usually easier to determine by off-the-shelf tools (Böcker & Dührkop, 2016; Goldman et al., 2023c). One insight of our work is to use those available tools by taking the chemical formula as given and developing a formula-constrained (i.e., heavy atom-constrained) generation pipeline. We also find it beneficial to exploit intermediate chemical representations to enable a scaled-up pretraining stage with theoretically unlimited fingerprint-structure pairs.
@@ 2.2 Inverse models for structure elucidation from spectra
However, the elucidation of structures that do not necessarily exist in any virtual chemical library requires generative techniques rather than retrieval-based techniques. MSNovelist (Stravs et al., 2022) builds an autoregressive fingerprint-to-molecule model that takes fingerprint predictions returned by CSI-FingerID and generates SMILES strings, with a decoding process that utilizes the molecular formula of the candidate compound. The molecular formula of an unknown molecule can be inferred from the MS1 and MS2 data with tools such as SIRIUS (Böcker & Dührkop, 2016) or MIST-CF (Goldman et al., 2023c). Spec2Mol (Litsa et al., 2023) develops a SMILES autoencoder and trains a spectrum CNN encoder model, with up to four spectral channels to accept spectra collected in low or high energy and positive or negative mode, that tries to predict the corresponding SMILES embedding from the spectrum. MassGenie (Shrivastava et al., 2021) is an orthogonal effort that uses forward MS models (Allen et al., 2015; Goldman et al., 2024) to augment training datasets with in silico reference spectra. Toward an end-to-end pipeline for molecular generation from mass spectra, which is the most relevant to our work, Butler et al. (2023) build MS2Mol, an end-to-end language model that encodes m/z values and intensities as tokenized text input and outputs an inferred chemical formula and SMILES string in an autoregressive manner. However, their implementation is not currently available at the time of writing, preventing direct comparison. Most recently, MADGEN (Wang et al., 2025) presents a diffusion generator of chemical structures from scaffolds as a two-stage generative process, seemingly bottlenecked in terms of accuracy by scaffold prediction. In this paper, we improve upon this thread of end-to-end approaches by encoding inductive biases via spectral transformers and utilizing a pretraining-finetuning framework for an MS-conditioned diffusion generator. DiffMS has two stages like MADGEN, but is trained end-to-end during its final training step, and is heavy-atom constrained.
@@ 2.3 Diffusion generative models
Denoising diffusion  (Sohl-Dickstein et al., 2015; Ho et al., 2020) has been shown to be widely effective across many tasks such as image (Song et al., 2020; Saharia et al., 2022; Karras et al., 2022) and text (Li et al., 2022; Austin et al., 2023) generation. More recently, diffusion has been applied to solve (bio)molecular generative tasks (Corso et al., 2023; Watson et al., 2023; Zeni et al., 2024).
@@ 3.3 Model parametrization and pretraining
We pretrain our diffusion decoder on a dataset of fingerprint-molecule pairs. Instead of using the spectrum encoder embeddings as the structural condition $\mathbf{y}$, we directly use the molecular fingerprint to condition the molecule generation. This is closely aligned with the mass spectra de novo generation task, as the decoder learns to generate molecules subject to strong structural constraints. Fingerprint-molecule datasets are essentially infinite in size, providing a promising path forward to further improve model performance by increasing the pretraining dataset size. To this end, we build a pretraining dataset consisting of 2.8M fingerprint-molecule pairs sampled from DSSTox (CCTE, 2019), HMDB (Wishart et al., 2021), COCONUT (Sorokina et al., 2021), and MOSES (Polykovskiy et al., 2020) datasets. Critically, we remove all NPLIB1 and MassSpecGym test and validation molecules from our decoder pretraining dataset so that our evaluation on the end-to-end generation task represents a setting where the model is generating truly novel structures. We provide an ablation and analysis of performance scaling with respect to pretraining dataset size in Sec. 4.4.
4 Experiments
4.1 Evaluation metrics
We adopt the de novo generation metrics from  Bushuiev et al. (2024):
@@ 4.1 Evaluation metrics
We report metrics for $k=1,10$ with additional figures in Appendix B. To obtain a ranked list of DiffMS predictions, we sample 100 molecules for each spectrum, remove invalid or disconnected molecules, and identify the top-$k$ molecules based on frequency. This post-processing is also applied to baseline methods for fairest comparison.
@@ 4.2 Datasets and baselines
MSNovelist (Stravs et al., 2022) builds a fingerprint-to-SMILES LSTM decoder to predict SMILES strings from the SIRIUS-generated CSI-FingerID fingerprint. The original implementation of MSNovelist is no longer functional, and furthermore relies on the closed-source CSI-FingerID fingerprint. The recently developed MIST model offers an open-source replacement with reported comparative performance to CSI-FingerID (Goldman et al., 2023b). Accordingly, we re-implement a baseline model that retains the main contributions of MSNovelist, adopting the code from Zhao et al. (2024), with a 4096-bit Morgan fingerprint spectral encoder using MIST alongside a formula-guided fingerprint-to-SMILES LSTM. This fingerprint-to-SMILES decoder is trained on the same 2.8M dataset used to pretrain our diffusion decoder; therefore, unlike the original MSNovelist implementation, both the spectral encoder and LSTM decoder never see any test structures. We use the ranking methodology from MSNovelist, wherein beam search (with a width of 100) and subsequently computed log-likelihoods are used for ranking. Similarly, Spec2Mol (Litsa et al., 2023) was also retrained on the NPLIB1 and MassSpecGym datasets for fair evaluation, with only one spectral channel instead of four used for training, to alleviate restrictions on collision energy or adduct. The same ranking for candidate molecules as used for DiffMS is applied.
@@ 4.3 Results
As seen in Table 1, DiffMS outperforms baseline methods on both datasets, including more than doubling the accuracy on MassSpecGym compared to the next best method, MADGEN. While there are several baseline methods that achieve non-zero prediction accuracy on NPLIB1, only MADGEN and DiffMS generate any correct structures on MassSpecGym. NPLIB1 is inherently a less challenging dataset than MassSpecGym; given the lack of a scaffold-based split, the CANOPUS test set contains many molecules that are nearly identical (Tanimoto similarity $>$ 0.85) to molecules in the train set (Bushuiev et al., 2024). This also explains the competitive performance of MIST + Neuraldecipher and MIST + MSNovelist, which both benefit from their ability to pretrain on these highly similar structures and learn to generate realistic structures as SMILES strings; the MSNovelist generation even more so given its formula-aware decoder. In contrast, MassSpecGym ensures that no molecules in the test set have an MCES $<$ 10 compared to any training molecule. As such, MassSpecGym evaluation represents a more challenging and more realistic out of distribution de novo generation setting, which illustrates the robust performance of DiffMS across all evaluation metrics. Examples of DiffMS-generated sampled are shown in Figure LABEL:fig:example_preds_main and Appendix D. In Appendix C, we show that even in cases where DiffMS does not recover the correct structure, it is consistently able to generate “close match“ structures that are still useful to domain experts.
@@ 4.4 Pretraining Ablations
Decoder Pretraining Ablation. We train DiffMS with increasing decoder pretraining dataset size, starting from 0 molecules (i.e., no pretraining) up to the full pretraining dataset of 2.8M molecules. As shown in Fig. 5, any amount of decoder pretraining offers a significant increase in performance. Additionally, we observe good performance scaling with increasing pretraining dataset size. Since fingerprint-to-molecule datasets are essentially infinite in size, this provides an avenue to continue to scale DiffMS’ performance by building even larger and more chemically comprehensive pretraining datasets.
Additional results and figures for encoder and decoder pretraining ablations can be found in Appendix B.
Figure 5: NPLIB1 top-$k$ accuracy for DiffMS pretrained on increasingly large fingerprint-to-molecule datasets. Full metrics available in Table 3 in the Appendix.
@@ Acknowledgments
This work was partly sponsored by DSO National Laboratories in Singapore, the MIT Summer Research Program (MSRP), National Science Foundation under grant IIS-2243850 (to S.J.), and ARPA-H under grant 1AY1AX000053 (to S.J.).
@@ Appendix A Experimental Details
We pretrain the encoder on the same dataset used for finetuning (i.e. NPLIB1, MassSpecGym), which are orders of magnitude smaller than the decoder pretraining dataset. For encoder pretraining, we use the multi-objective loss settings of Goldman et al. (2023b). We pretrain the encoder for 100 epochs using the RAdam optimizer  (Liu et al., 2021) and a constant learning rate.
We finetune the end-to-end model using cross-entropy loss and no auxiliary training objectives, i.e. only the denoising diffusion objective. We use the AdamW optimizer with cosine annealing learning rate schedule for finetuning. We find that finetuning on NPLIB1 requires more epochs because the dataset is much smaller than MassSpecGym. As such, on NPLIB1, we finetune for 250 epochs and on MassSpecGym, we finetune for 50 epochs.
Appendix B Additional Results on Ablations
We provide additional results for ablations studies in Table 3, Fig LABEL:figure:appendix_ablate_encoder, and Fig. LABEL:figure:appendix_ablate_decoder.
@@ Appendix B Additional Results on Ablations
2.8M 8.34% 11.95 0.35 15.19% 9.23 0.46
Appendix C Evaluations and baselines
As an addendum to the evaluations in Table 1, we provide an comparison of additional metrics for additional context of performance. Namely, we evaluate the percentage of valid molecules generated by the method and the percentage of molecules that have a formula match to the true molecule. Adopting the domain-expert thresholds put forth by MS2Mol (Butler et al., 2023), we also evaluate whether candidate molecules were a “meaningful” match in structural similarity, having a Tanimoto similarity of 0.4 or greater; or a “close match” in structural similarity, having a Tanimoto similarity of 0.675 or greater. These statistics are reported in Table 4.
Table 4: Additional evaluation of molecule validity, formula matching, and appearance above domain-expert-defined Tanimoto thresholds on de novo structural elucidation performance on MassSpecGym (Bushuiev et al., 2024). The best performing model for each metric is bold and the second best is underlined. * indicates our implementations of baseline approaches. Definitions of meaningful match (Tanimoto similarity $\geq 0.4$) and close match (Tanimoto similarity $\geq 0.675$) are taken from (Butler et al., 2023).
Overall Top-1 Top-10
Model % Valid $\uparrow$ % Meaningful match $\uparrow$ % Close match $\uparrow$ % Meaningful match $\uparrow$ % Close match $\uparrow$
NPLIB1
Spec2Mol 66.5% 0.00% 0.00% 0.00% 0.00%
MIST + Neuraldecipher 91.11% 29.30% 7.33% 41.39% 12.82%
MIST + MSNovelist 98.60% 32.90% 11.78% 44.79% 19.02%
DiffMS 100.0% 27.40% 12.83% 46.45% 22.04%
MassSpecGym
Spec2Mol 68.5% 0.0% 0.0% 0.0% 0.0%
MIST + Neuraldecipher 81.78% 0.29% 0.01% 0.39% 0.09%
MIST + MSNovelist 98.58% 0.66% 0.00% 1.92% 0.00%
DiffMS 100.0% 12.41% 3.78% 32.47% 6.73%
Appendix D DiffMS Generated Molecules
D.1 NPLIB1 Molecules
D.2 MassSpecGym Molecules

### v2

@@ 1 Introduction
In this paper, we focus on the “inverse” MS problem and develop a novel machine learning framework for chemical structure generation given a mass spectrum, sometimes also described as de novo generation. A recent study of de novo generation from MS shows that all of the methods tested suffer from near-zero structural accuracy (Bushuiev et al., 2024). Among prior approaches to this task are language models that are trained to convert tokenized m/z values and intensities as inputs to SMILES strings as outputs (Butler et al., 2023; Litsa et al., 2023); however, these autoregressive language models do not capture the permutation-invariant nature of mass spectra and molecules, nor do they incorporate chemical formula constraints as helpful prior knowledge. Another family of approaches utilizes intermediate representations such as scaffolds (Wang et al., 2025) or fingerprints (Stravs et al., 2022) before generating chemical structures, which are arguably more chemically interpretable and leverage additional amounts of structure-only data available, but have not necessarily led to significant performance improvement on benchmarks. Compared to the complete structural elucidation challenge of the “inverse” MS problem, the chemical formula of the unknown molecule is usually easier to determine by off-the-shelf tools from MS1 and MS2 data, utilizing tools such as SIRIUS (Böcker & Dührkop, 2016), BUDDY (Xing et al., 2023), or MIST-CF (Goldman et al., 2023c). One insight of our work is to use those available tools by taking the chemical formula as given and developing a formula-constrained (i.e., heavy atom-constrained) generation pipeline. We also find it beneficial to exploit intermediate chemical representations to enable a scaled-up pretraining stage with theoretically unlimited fingerprint-structure pairs.
@@ 2.2 Inverse models for structure elucidation from spectra
However, the elucidation of structures that do not necessarily exist in any virtual chemical library requires generative techniques rather than retrieval-based techniques. MSNovelist (Stravs et al., 2022) builds an autoregressive fingerprint-to-molecule model that takes fingerprint predictions returned by CSI-FingerID and generates SMILES strings, with a decoding process that utilizes the molecular formula of the candidate compound as inferred from tools such as SIRIUS(Böcker & Dührkop, 2016) or MIST-CF (Goldman et al., 2023c). Spec2Mol (Litsa et al., 2023) develops a SMILES autoencoder and trains a spectrum CNN encoder model, with up to four spectral channels to accept spectra collected in low or high energy and positive or negative mode, that tries to predict the corresponding SMILES embedding from the spectrum. MassGenie (Shrivastava et al., 2021) is an orthogonal effort that uses forward MS models (Allen et al., 2015; Goldman et al., 2024) to augment training datasets with in silico reference spectra. Toward an end-to-end pipeline for molecular generation from mass spectra, which is the most relevant to our work, Butler et al. (2023) build MS2Mol, an end-to-end language model that encodes m/z values and intensities as tokenized text input and outputs an inferred chemical formula and SMILES string in an autoregressive manner. However, their implementation is not currently available at the time of writing, preventing direct comparison. Most recently, MADGEN (Wang et al., 2025) presents a diffusion generator of chemical structures from scaffolds as a two-stage generative process, seemingly bottlenecked in terms of accuracy by scaffold prediction. In this paper, we improve upon this thread of end-to-end approaches by encoding inductive biases via spectral transformers and utilizing a pretraining-finetuning framework for an MS-conditioned diffusion generator. DiffMS has two stages like MADGEN, but is trained end-to-end during its final training step, and is heavy-atom constrained.
@@ 2.3 Diffusion generative models
Denoising diffusion  (Sohl-Dickstein et al., 2015; Ho et al., 2020) has been shown to be widely effective across many tasks such as image (Song et al., 2021; Saharia et al., 2022; Karras et al., 2022) and text (Li et al., 2022; Austin et al., 2023) generation. More recently, diffusion has been applied to solve (bio)molecular generative tasks (Corso et al., 2023; Watson et al., 2023; Zeni et al., 2024).
@@ 3.3 Model parametrization and pretraining
We pretrain our diffusion decoder on a dataset of fingerprint-molecule pairs. Instead of using the spectrum encoder embeddings as the structural condition $\mathbf{y}$, we directly use the molecular fingerprint to condition the molecule generation. This is closely aligned with the mass spectra de novo generation task, as the decoder learns to generate molecules subject to strong structural constraints. Fingerprint-molecule datasets are essentially infinite in size, providing a promising path forward to further improve model performance by increasing the pretraining dataset size. To this end, we build a pretraining dataset consisting of 2.8M fingerprint-molecule pairs sampled from DSSTox (CCTE, 2019), HMDB (Wishart et al., 2021), COCONUT (Sorokina et al., 2021), and MOSES (Polykovskiy et al., 2020) datasets. Critically, we remove all NPLIB1 and MassSpecGym test and validation molecules from our decoder pretraining dataset so that our evaluation on the end-to-end generation task represents a setting where the model is generating truly novel structures. Bushuiev et al. (2024) provide their own dataset of 4M molecules, but use a different exclusion criteria to prevent data leakage. We provide an ablation and analysis of performance scaling with respect to pretraining dataset size in Sec. 4.4.
@@
4 Experiments
4.1 Evaluation metrics
We adopt the de novo generation metrics from  Bushuiev et al. (2024):
We report metrics for $k=1,10$ with additional results in Appendix C. To obtain a ranked list of DiffMS predictions, we sample 100 molecules for each spectrum, remove invalid or disconnected molecules, and identify the top-$k$ molecules based on frequency. This post-processing is also applied to baseline methods for fairest comparison.
@@ 4.2 Datasets and baselines
MSNovelist (Stravs et al., 2022) builds a fingerprint-to-SMILES LSTM decoder to predict SMILES strings from the SIRIUS-generated CSI-FingerID fingerprint. The original implementation of MSNovelist is not readily retrainable, and furthermore relies on the closed-source CSI-FingerID fingerprint. The recently developed MIST model offers an open-source replacement with reported comparative performance to CSI-FingerID (Goldman et al., 2023b). Accordingly, we re-implement a baseline model that retains the main contributions of MSNovelist, adopting the code from Zhao et al. (2024), with a 4096-bit Morgan fingerprint spectral encoder using MIST alongside a formula-guided fingerprint-to-SMILES LSTM. This fingerprint-to-SMILES decoder is trained on the same 2.8M dataset used to pretrain our diffusion decoder; therefore, unlike the original MSNovelist implementation, both the spectral encoder and LSTM decoder never see any test structures. We use the ranking methodology from MSNovelist, wherein beam search (with a width of 100) and subsequently computed log-likelihoods are used for ranking. Similarly, Spec2Mol (Litsa et al., 2023) was also retrained on the NPLIB1 and MassSpecGym datasets for fair evaluation, with only one spectral channel instead of four used for training, to alleviate restrictions on collision energy or adduct. The same ranking for candidate molecules as used for DiffMS is applied.
@@ 4.3 Results
As seen in Table 1, DiffMS outperforms baseline methods on both datasets, including more than doubling the accuracy on MassSpecGym compared to the next best method, MADGEN. While there are several baseline methods that achieve non-zero prediction accuracy on NPLIB1, only MADGEN and DiffMS generate any correct structures on MassSpecGym. NPLIB1 is inherently a less challenging dataset than MassSpecGym; given the lack of a scaffold-based split, the CANOPUS test set contains many molecules that are nearly identical (Tanimoto similarity $>$ 0.85) to molecules in the train set (Bushuiev et al., 2024). This also explains the competitive performance of MIST + Neuraldecipher and MIST + MSNovelist, which both benefit from their ability to pretrain on these highly similar structures and learn to generate realistic structures as SMILES strings; the MSNovelist generation even more so given its formula-aware decoder. In contrast, MassSpecGym ensures that no molecules in the test set have an MCES $<$ 10 compared to any training molecule. As such, MassSpecGym evaluation represents a more challenging and more realistic out of distribution de novo generation setting, which illustrates the robust performance of DiffMS across all evaluation metrics. Examples of DiffMS-generated sampled are shown in Figure LABEL:fig:example_preds_main and Appendix E. In Appendix B, we show that even in cases where DiffMS does not recover the correct structure, it is consistently able to generate “close match“ structures that are still useful to domain experts.
@@ 4.4 Pretraining Ablations
Table 3: DiffMS de novo structural elucidation performance on NPLIB1 (Dührkop et al., 2021b) and MassSpecGym (Bushuiev et al., 2024) datasets using MIST-CF annotated formulae and ground truth formulae. The best performing model for each metric is bold.
Top-1 Top-10
Formulae Accuracy $\uparrow$ MCES $\downarrow$ Tanimoto $\uparrow$ Accuracy $\uparrow$ MCES $\downarrow$ Tanimoto $\uparrow$
NPLIB1
MIST-CF Formulae 7.03% 11.81 0.36 14.98% 9.39 0.48
True Formulae 8.34% 11.95 0.35 15.44% 9.23 0.47
MassSpecGym
MIST-CF Formulae 1.86% 17.83 0.27 4.10% 13.71 0.40
True Formulae 2.30% 18.45 0.28 4.25% 14.73 0.39
@@
Figure 7: NPLIB1 top-$k$ accuracy for DiffMS pretrained on increasingly large fingerprint-to-molecule datasets. Additional metrics available in Table 5 in the Appendix.
Decoder Pretraining Ablation. We train DiffMS with increasing decoder pretraining dataset size, starting from 0 molecules (i.e., no pretraining) up to the full pretraining dataset of 2.8M molecules. As shown in Fig. 7, any amount of decoder pretraining offers a significant increase in performance. Additionally, we observe good performance scaling with increasing pretraining dataset size. Since fingerprint-to-molecule datasets are essentially infinite in size, this provides an avenue to continue scaling DiffMS’ performance by building even larger and more chemically comprehensive pretraining datasets.
Additional results and figures for encoder and decoder pretraining ablations can be found in Appendix C.
4.5 Formula Inference Ablation
In many real-world elucidation settings, chemists may know the true chemical formula of the target compound a priori or be able to determine the true formula using auxiliary methods. However, chemical formulae can also be predicted from the spectrum with high accuracy by out-of-the-box formula annotation tools (Goldman et al., 2023c; Xing et al., 2023; Böcker & Dührkop, 2016). In this section, we broaden the structural elucidation challenge and investigate the ability of DiffMS to rely on MIST-CF (Goldman et al., 2023c) formula predictions to test its performance in settings where the true chemical formula is unknown.
For each spectrum, we predict the top 5 most likely formulae using MIST-CF. We then generate candidate structures for each of the predicted formulae. To have a fair comparison, we still generate 100 total molecules, split across the 5 predicted formulae. As shown in Table 3, DiffMS still has strong performance even when relying on formula annotation tools to supply the formula. While the elucidation accuracy is slightly lower, the MCES and Tanimoto metrics are actually better in some cases using the MIST-CF predicted formulae. Intuitively, sampling molecules with different formulae gives us higher diversity and thus a better chance of getting a “close” structure.
@@ Acknowledgments
This work was partly sponsored by DSO National Laboratories in Singapore (to C.W.C.), the MIT Summer Research Program (MSRP), National Science Foundation under grant IIS-2243850 (to S.J.), and ARPA-H under grant 1AY1AX000053 (to S.J.).
@@ Appendix A Experimental Details
We pretrain the encoder on the same dataset used for finetuning (i.e. NPLIB1, MassSpecGym), which are orders of magnitude smaller than the decoder pretraining dataset. For encoder pretraining, we use the multi-objective loss settings of Goldman et al. (2023b). We pretrain the encoder for 100 epochs using the RAdam optimizer  (Liu et al., 2020).
We finetune the end-to-end model using cross-entropy loss and no auxiliary training objectives, i.e. only the denoising diffusion objective. We use the AdamW optimizer with cosine annealing learning rate schedule for finetuning. We finetune DiffMS for 50 epochs on NPLIB1 and 15 epochs on MassSpecGym.
DiffMS is a relatively lightweight model, and all experiments were run on NVIDIA 2080ti GPUs with 12 GB of memory. On these GPUs, finetuning DiffMS takes 1.45 minutes per epoch on CANOPUS and 46 minutes per epoch on MassSpecGym. It takes 4 minutes on average to generate 100 samples from DiffMS.
Appendix B Additional Results
As an addendum to the evaluations in Table 1, we provide some additional metrics to further contextualize DiffMS performance. Firstly, we evaluate the percentage of model samples that correspond to valid molecules. Additionally, we adopt the domain-expert thresholds put forth by MS2Mol (Butler et al., 2023), where we evaluate whether candidate molecules were a “meaningful” match in structural similarity, having a Tanimoto similarity of 0.4 or greater; or a “close match” in structural similarity, having a Tanimoto similarity of 0.675 or greater. We omit MADGEN and the baseline methods from Bushuiev et al. (2024) as they do not report these metrics.
As shown in Table 4, 100% of DiffMS samples are valid molecules. This is directly enforced because of our graph-based representation. In contrast, SMILES strings generated by baseline methods may not correspond to a valid structure. We find that DiffMS consistently achieves higher meaningful and close match rates than baseline methods. Impressively, DiffMS achieves over 32 times more meaningful matches in the top-10 predictions than the next best baseline on MassSpecGym. These results show that while generating exact matches continues to be a challenging task for de novo structural elucidation, DiffMS is able to generate meaningful structural matches at a high rate.
@@ Appendix B Additional Results
2.8M 8.34% 11.95 0.35 15.44% 9.23 0.47
Appendix C Ablations
C.1 Additional Pretraining Ablation Results
In this section, we provide additional results for the ablation studies in Sec. 4.4. Table 5 and Fig. LABEL:figure:appendix_ablate_decoder demonstrate DiffMS’ performance scaling with respect to increasingly large decoder pretraining datasets, and Fig LABEL:figure:appendix_ablate_encoder shows the impact of pretraining the spectra encoder.
C.2 Prior Distribution Ablations
In this section we provide an additional ablation study to justify the choice of the marginal prior distribution. Specifically, we compare with two alternative prior distributions: the “empty” distribution, consisting of no bonds, and the “fully connected” distribution, consisting of all single bonds. As shown in Table 6, the marginal distribution performs best, though the empty distribution is not far behind. Intuitively, the empty distribution is close to the marginal distribution as molecular graphs are typically very sparse. These results support our intuitions that having a prior distribution that is closer to the data distribution results in better performance.
Table 6: DiffMS performance on NPLIB1 with different prior distributions. The best performing model for each metric is bold and second best is underlined.
Prior Distribution Top-1 Top-10
Fully Connected 3.36% 12.67 0.28 7.60% 9.56 0.4
Empty 6.60% 11.55 0.34 14.94% 9.07 0.47
Marginal 8.34% 11.95 0.35 15.44% 9.23 0.47
Appendix D Formulae Annotation Study
In this section, we provide additional experiments using MIST-CF (Goldman et al., 2023b) and BUDDY (Xing et al., 2023) for formula annotation on the NPLIB1 and MassSpecGym datasets. Specifically, we use BUDDY and MIST-CF to predict the top-5 most likely formulae for each spectra in the test sets and measure the accuracy of these formula annotations.
As shown in Table 7, MIST-CF and BUDDY both achieve good performance on NPLIB1, where MIST-CF achieves over 90% top-5 accuracy. However, both methods struggle on MassSpecGym, underscoring the difficulty of this dataset. It is important to note that neither NPLIB1 nor MassSpecGym include MS1 data, such as precursor m/z, which can aid in deriving accurate formula annotations. As such, these formula annotation accuracies are likely lower than what could be achieved in end-to-end elucidation workflows.
Table 7: Formula annotation accuracy for MIST-CF (Goldman et al., 2023c) and BUDDY(Xing et al., 2023) on the NPLIB1 (Böcker & Dührkop, 2016) and MassSpecGym (Bushuiev et al., 2024) datasets. The best performing method for each metric is bold.
Model NPLIB1 MassSpecGym
Top-1 Acc. Top-5 Acc. Top-1 Acc. Top-5 Acc.
BUDDY 78% 83% 59% 71%
MIST-CF 84% 92% 48% 69%
Appendix E DiffMS Generated Molecules
E.1 NPLIB1 Molecules
E.2 MassSpecGym Molecules